• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name gilman

    Log entry time 08:04:54 on October 9,1999

    Entry number 28033

    keyword=oct 9 owl shift summary

    Start shift finishing access to check FPP ch 1 U problem.
    Somewhat obscure, probably loose cable, it gets fixed
    by X JIang and R Gilman.

    00:55 - 1:05: finished 1H radiator out, last run 2173,
    switch to 2H target

    1:05 - 2H radiator out runs, first run 2174
    2:30 beam down for a while for tuning
    3:00 beam stability starts stinking
    3:30 done with run 2176, time to move radiator in, ask
    for current reduction to 30 microamps.

    3:30 move radiator in, found previous person to move it out left it enabled.
    Finger breaking lessons needed?

    3:40 the hour of stinky beam continues difficult to take a run with trips
    and beam offcenter.

    run 2177 garbage, run 2178 first gamma + d run.

    4:00 Hall C shuts down for hour, waiting for injector person, as MCC
    cannot keep beam up at 60 microamps. We are getting stable
    beam in meantime.

    5:08 beam stable at 30.5 uA with no trips for 60 + minutes, now off for
    injector expert investigation, an hour or so

    6:05 MCC calls and says they are about to start restoring beam

    6:50 As MCC hasbeen playing with beam for a while, figuring taking data
    while they play cannot hurt. Run 2181 crashes, start 2182.

    7:30 MCC says they are done playing with beam, coda does a big crash, TS1
    7:40 start run 2182 TS1 crashes to end it